OpenStudy (anonymous):

What is the fifth term of a sequence whose first term is 5 and whose common ratio is 3? a) 243 b) 405 c) 1,215

OpenStudy (anonymous):

??????????

OpenStudy (sgadi):

5*3^4=81*5=405